# Polynomial entropy on the $n$-fold symmetric product and its suspension

### v1

@@ Abstract
We prove that the polynomial entropy of the induced map $F_{n}(f)$ on the $n$-fold symmetric product of a compact space $X$ and its suspension are both equal to $nh_{pol}(f)$, when $f:X\to X$ is a homeomorphism with a finite non-wandering set $NW(f)$. We also give a lower bound for the polynomial entropy on the suspension, for a homeomorphism $f$ with at least one wandering point, under certain assumptions.
@@ 1 Introduction
In [12], the authors computed the polynomial entropy of the induced maps $f^{*k}$ and $C(f)$ when $f$ is a homeomorphism of a circle or an interval with finitely many non-wandering points. Those results heavily rely on the coding techniques developed in [15] and slightly generalized in [17], which apply only to the case when the non-wandering set is finite (so every non-wandering point is actually periodic).
In the present paper, we prove that the polynomial entropy of the dynamical system $(\mathcal{S}F_{n}(X),\mathcal{S}F_{n}(f))$ is equal to the polynomial entropy of $(F_{n}(X),F_{n}(f))$, and that they are both equal $nh_{pol}(f,X)$, under the assumptions that $X$ is compact and $f:X\to X$ is a homeomorphism with a finite non-wandering set. The condition regarding the finite number of non-wandering points comes naturally, because of the aforementioned growth of wandering orbits, which is polynomial. Comparing to the results obtained in [12] for $(F_{n}(X),F_{n}(f))$, we generalise them for an arbitrary compact $X$ and complement them with results regarding the suspension (and subsequently the generalised suspension).
@@
Let $X$ be a compact space, $f:X\to X$ a homeomorphism with a finite non-wandering set and $n\geqslant 2$. Then
@@
Furthermore, in contrast to the condition of a finite non-wandering set, we turn our attention to homeomorphisms with at least one wandering point. We obtain a similar result to that in [13]:
@@ 2.1 Polynomial entropy and coding
If $X=Y$, we abbreviate $h_{pol}(f):=h_{pol}(f;X)$. The polynomial entropy, as well as the topological entropy, can also be defined via coverings with sets of $d^{n}_{f}$-diameters less than $\varepsilon$, or via coverings by balls of $d_{n}^{f}$-radius less than $\varepsilon$, see p. $626$ in [26]. We list some properties of the polynomial entropy that are important for our computations (for the proofs see Propositions $1-4$ in [26]).
@@
We abbreviate $h_{pol}(\mathcal{F};Y):=h_{pol}(f,\mathcal{F};Y)$ whenever there is no risk of confusion. When we are dealing with a noncompact $Y$, we can relate the polynomial entropy $h_{pol}(f;Y)$ and the polynomial entropy relative to a one point family =$\mathcal{F}=\{K\cap Y\}$, for all compact $K\subseteq X$:
@@ 2.2 Hyperspaces and induced maps
If $f$ is a homeomorphism, then this is also true for both $2^{f}$ and $F_{n}(f)$. Because of the following commutative diagram, it is clear that $(F_{n}(f),F_{n}(X))$ is a factor of $(X^{\times n},f^{\times n})$, so $h_{pol}(F_{n}(f),F_{n}(X))\leqslant h_{pol}(X^{\times n},f^{\times n})=nh_{pol}(f,X)$.
[…]
@@ 3 Main results and proofs
Let $f:X\to X$ be a homeomorphism with finitely many non-wandering points. Since $NW(f)$ is finite and invariant, we conclude it consists of only periodic points. We can pass from $f$ to $f^{k}$, for some $k\geqslant 1$, so that all the non-wandering points become fixed. This change does not affect the polynomial entropy, since we have that $h_{pol}(f^{k})=h_{pol}(f)$. It is clear that this is also true for the map $F_{n}(f)$, i.e.
@@
=$=\left\{\begin{array}[]{ll}qF_{n}(f)^{k}q^{-1}(\mathcal{A})&\text{if }\mathcal{A}\neq F_{X}\\ F_{X}&\text{if }\mathcal{A}=F_{X}\end{array}\right.$
@@
so we can, from now on, consider homeomorphisms $f$ with $NW(f)=Fix(f)$.
@@
Let $X$ be a compact space and $f:X\to X$ a homeomorphism with a finite non-wandering set $NW(f)=Fix(f)$. Then $\mathcal{S}F_{n}(f))$ has a finite non-wandering set and $NW(\mathcal{S}F_{n}(f)))=Fix(\mathcal{S}F_{n}(f)))$, for all $n\geqslant 2$.
@@
Since $U$ was an arbitrary neighbourhood of $A$, we see that $A\in NW(F_{n}(X))$. From Proposition 5 in [12] we have that $NW(F_{n}(X))=Fix(F_{n}(X))$, and finally:
@@
Let $X$ be a compact space, $f:X\to X$ a homeomorphism with a finite non-wandering set and $n\geqslant 2$. Then
@@
As we explained, we can consider all the non-wandering points to be fixed points. Let
@@
First note that $\pi^{-1}(G_{n})$ is $f^{\times n}$-invariant, since $f$ and the restriction of $\pi$ are bijections. Following the same reasoning as before, there is either exactly one non-wandering point, or there is none (either instantaneously, or after removing all the fixed points), so we add one point to the sets $\pi^{-1}(G_{n})$ and $G_{n}$ point $(\tilde{x},\ldots,\tilde{x})$ and $\{\tilde{x}\}$, respectively. $H$ will remain a homeomorphism, and the entropies are not changed. We can conclude that
@@
Let $X$ be a compact space, $f:X\to X$ a homeomorphism with a finite non-wandering set and $n>m\geqslant 1$. Then
@@
As we mentioned in the introduction, we obtained the results regarding the polynomial entropy on $F_{n}(X)$, for specific spaces $X=[0,1]$ and $\mathbb{S}^{1}$ in [12]. We complement them with the results on the symmetric product suspension, using the main Theorem 7 and the fact that $h_{pol}(f,[0,1])=h_{pol}(f,\mathbb{S}^{1})=1$ when the set $NW(f)$ is finite.
@@
We will use the same notation as in the proof of Theorem 6 in [13]. Define $x_{n}=(f(x_{0}))$, $n\in\mathbb{Z}$ and
@@
We can assume that $f$ is strictly increasing and that $f\neq\mathrm{Id}$ (otherwise, consider $f^{2}$ instead of $f$). Since in this case ${Fix}(f)={NW}(f)$, there exists at least one wandering point $x_{0}$. In order to deduce this result we only have to comment on the fact that $\alpha_{f}(x_{0})$ and $\omega_{f}(x_{0})$ are in fact one-element sets, since $\lim\limits_{n\to\pm\infty}f^{n}(x_{0})$ exists. ∎

### v2

@@ Abstract
We prove that the polynomial entropy of the induced map $F_{n}(f)$ on the $n$-fold symmetric product of a compact space $X$ and its suspension are both equal to $nh_{pol}(f)$, when $f:X\to X$ is a homeomorphism with a finite chain recurrent set $\mathcal{CR}(f)$. We also give a lower bound for the polynomial entropy on the suspension, for a homeomorphism $f$ with at least one wandering point, under certain assumptions.
@@ 1 Introduction
In [12], the authors computed the polynomial entropy of the induced maps $F_{n}(f)$ and $C(f)$ when $f$ is a homeomorphism of a circle or an interval with finitely many non-wandering points. Those results heavily rely on the coding techniques developed in [15] and slightly generalized in [17], which apply only to the case when the non-wandering set is finite (so every non-wandering point is actually periodic). The condition regarding the finite number of non-wandering points comes naturally, because of the aforementioned growth of wandering orbits, which is polynomial.
In the present paper, we prove that the polynomial entropy of the dynamical system $(\mathcal{S}F_{n}(X),\mathcal{S}F_{n}(f))$ is equal to the polynomial entropy of $(F_{n}(X),F_{n}(f))$, and that they are both equal $nh_{pol}(f,X)$, under the assumptions that $X$ is compact and $f:X\to X$ is a homeomorphism with a finite chain recurrent set. Since it is needed that $F_{n}(f)$ has a finite non-wandering set in order to apply the aforementioned method, the assumptions for map $f$ have to be a little stronger than $NW(f)$ being finite. One sufficient condition is that $\mathcal{CR}(f)$ is finite. If $X$ is, for example, a regular curve, we have that actually $\mathcal{CR}(f)=NW(f)$ (see [28]). It is worth mentioning that dendrites and graphs are regular curves.
Comparing to the results obtained in [12] for $(F_{n}(f),F_{n}(X))$, we generalise them for an arbitrary homeomorphism on a compact $X$ with a finite chain recurrent set and complement them with results regarding the suspension (and subsequently the generalised suspension).
@@
Let $X$ be a compact space, $f:X\to X$ a homeomorphism with a finite chain recurrent set and $n\geqslant 2$. Then
@@
Furthermore, in contrast to the condition of a finite chain recurrent set, we turn our attention to homeomorphisms with at least one wandering point. We obtain a similar result to that in [13]:
@@ 2.1 Polynomial entropy and coding
If $X=Y$, we abbreviate $h_{pol}(f):=h_{pol}(f;X)$. The polynomial entropy, as well as the topological entropy, can also be defined via coverings with sets of $d^{n}_{f}$-diameters less than $\varepsilon$, or via coverings by balls of $d_{n}^{f}$-radius less than $\varepsilon$, see p. $626$ in [26]. We list some properties of the polynomial entropy that are important for our computations (for the proofs see Propositions 141-4 in [26]).
@@
We abbreviate $h_{pol}(\mathcal{F};Y):=h_{pol}(f,\mathcal{F};Y)$ whenever there is no risk of confusion. When we are dealing with a noncompact $Y$, we can relate the polynomial entropy $h_{pol}(f;Y)$ and the polynomial entropy relative to a one point family $\mathcal{F}=\{K\cap Y\}$, for all compact $K\subseteq X$:
@@ 2.2 Hyperspaces and induced maps
If $f$ is a homeomorphism, then this is also true for both $2^{f}$ and $F_{n}(f)$. Because of the following commutative diagram, it is clear that $(F_{n}(f),F_{n}(X))$ is a factor of $(f^{\times n},X^{\times n})$, so $h_{pol}(F_{n}(f),F_{n}(X))\leqslant h_{pol}(f^{\times n},X^{\times n})=nh_{pol}(f,X)$.
$X^{\times n}$$X^{\times n}$$F_{n}(X)$$F_{n}(X)$$\scriptstyle{f^{\times n}}$$\scriptstyle{\pi}$$\scriptstyle{\pi}$$\scriptstyle{F_{n}(f)}$
@@ 2.3 Suspensions
3 Finiteness of $NW(F_{n}(f))$
As we already mentioned, in order to apply the method described in [15] and in [17], the non-wandering set of $F_{n}(f)$ has to be finite. As we can see from Examples 14 and 15 in [31], the non-wandering set of the induced map $F_{n}(f)$ is not directly connected with the non-wandering set of $f$, as one might think. One way to achieve finiteness of $NW(F_{n}(f))$ is that we assume that $\mathcal{CR}(f)$ is finite, as we will now prove. Let us start with an auxiliary lemma.
Lemma 6.
Let $X$ be a compact metric space and $f:X\to X$ a continuous map. Let $A,B\in F_{N}(X)$, $A=\{a_{1},\ldots,a_{k}\}$ and $B=\{b_{1},\ldots,b_{m}\}$, $1\leqslant m,k\leqslant N$. There exists $\varepsilon_{0}$ such that for every $\varepsilon<\varepsilon_{0}$ the following is true: if $d_{H}(A,B)<\varepsilon$, then there exist NN-tuples $(\alpha_{1},\ldots,\alpha_{N})\in\pi^{-1}(A)$ and $(\beta_{1},\ldots,\beta_{N})\in\pi^{-1}(B)$ such that $d((\alpha_{1},\ldots,\alpha_{N}),(\beta_{1},\ldots,\beta_{N}))<\varepsilon$.
Proof.
Let $\delta:=\min\limits_{i\neq j}{d(a_{i},a_{j})}$, $\varepsilon_{0}:=\delta/2$, $\varepsilon<\varepsilon_{0}$ and suppose that $d_{H}(A,B)<\varepsilon$. Then, from the definition of Hausdorff metric, we have that
$(\forall i\in\{1,\ldots,k\})(\exists j_{i}\in\{1,\ldots,m\})\ d(a_{i},b_{j_{i}})<\varepsilon.$
Note that for $i\neq l$, we have that $b_{j_{i}}\neq b_{j_{l}}$. Indeed,
$\delta\leqslant d(a_{i},a_{l})$ $\leqslant d(a_{i},b_{j_{i}})+d(b_{j_{i},b_{j_{l}}})+d(b_{j_{l}},a_{l})$
$<2\varepsilon+d(b_{j_{i}},b_{j_{l}})<\delta+d(b_{j_{i}},b_{j_{l}}).$
Therefore $d(b_{j_{i}},b_{j_{l}})>0$, and consequently $b_{j_{i}}\neq b_{j_{l}}$. We conclude that it has to be $m\geqslant k$.
We can now explain how we define $(\alpha_{1},\ldots,\alpha_{N})\in\pi^{-1}(A)$ and $(\beta_{1},\ldots,\beta_{N})\in\pi^{-1}(B)$. Firstly, $(\alpha_{1},\ldots,\alpha_{k})=(a_{1},\ldots,a_{k})$ and $(\beta_{1},\ldots,\beta_{k})=(b_{j_{1}},\ldots,b_{j_{k}})$. Then let $\{\beta_{j_{k}+1},\ldots,\beta_{m}\}$ be the elements from $\{b_{1},\ldots,b_{m}\}\setminus\{b_{j_{1}},\ldots,b_{j_{k}}\}$. The corresponding $\{\alpha_{j_{k}+1},\ldots,\alpha_{m}\}$ can be chosen from $\{a_{1},\ldots,a_{k}\}$, in such way that $d(\alpha_{p},\beta_{p})<\varepsilon$, for $p\in\{j_{k}+1,\ldots,m\}$. Indeed, if this weren’t the case, then $d_{H}(A,B)\geqslant\varepsilon$. For the remainder of the pairs of coordinates, if there are any ($m<N$), we can just repeat any of the already used pairs. In this way we obtained $N$-touples which are close, when their projections in $F_{N}(X)$ are close.
Theorem 7.
Let $X$ be a compact metric space and $f:X\to X$ a continuous map such that $\mathcal{CR}(f)=\mathrm{Fix}(f)=\mathrm{NW}(f)$ is a finite set. Then
$\mathrm{NW}(F_{N})(f)=\mathrm{Fix}(F_{N})=\big\{\{x_{1},\ldots,x_{k}\}\mid 1\leq k\leq N,\,x_{j}\in\mathrm{Fix}(f)\big\}.$
Proof.
Let $A\in\mathrm{NW}(F_{N})(f)$, $A=\{a_{1},\ldots,a_{k}\}$, $k\leqslant N$. If we show that $(a_{1},\ldots,a_{k})\in\mathcal{CR}(f^{\times k})$, $f^{\times k}:X^{\times k}\to X^{\times k}$, then using the fact that $\mathcal{CR}(f\times g)=\mathcal{CR}(f)\times\mathcal{CR}(g)$, we have that $a_{i}\in\mathcal{CR}(f)$, $1\leqslant i\leqslant k$. Since $\mathcal{CR}(f)=\mathrm{Fix}(f)$, we have that $A\in\mathrm{Fix}(F_{N}(f))$. Let us show that $(a_{1},\ldots,a_{k})\in\mathcal{CR}(f^{\times k})$. More precisely, we will construct an $\varepsilon$-chain starting and ending at $(a_{1},\ldots,a_{k})$, for an arbitrary $\varepsilon>0$. Note that it is enough to find an $\varepsilon$-chain starting and ending at any permutation of $(a_{1},\ldots,a_{k})$.
Since is $A$ is non-wandering, for every $n\in\mathbb{N}$, choose $U_{n}:=B(A,1/n)$. There exists $B_{n}\in F_{N}(X)$ and $m_{n}\in\mathbb{N}$ such that
$d(A,B_{n})<1/n,\quad d(A,F_{N}^{m_{n}}(f)(B_{n}))<1/n.$
We can also choose $(m_{n})_{n\in\mathbb{N}}$ to be strictly increasing sequence.
Denote by $\pi^{-1}(A)=\{\mathbf{a}_{1},\ldots,\mathbf{a}_{p}\}\in X^{N}$, $p\leqslant P$ and by $\pi^{-1}(B_{n})=\{\mathbf{b}_{1}^{n},\ldots,\mathbf{b}_{q(n)}^{n}\}\in X^{N}$, $q(n)\leqslant Q$. Using Lemma 6, we can conclude that for every $n\in\mathbb{N}$, there exist $i_{n},j_{n}\in\{1,\ldots,p\}$ and $l_{n},r_{n}\in\{1,\ldots,q(n)\}$ such that
$d(\mathbf{a}_{i_{n}},\mathbf{b}_{r_{n}}^{n})<1/n,\quad d\left(\mathbf{a}_{j_{n}},{(f^{\times N})}^{m_{n}}(\mathbf{b}_{l_{n}}^{n})\right)<1/n.$
Since the set $\{1,\ldots,p\}$ is finite, there exists $i_{0}\in\{1,\ldots,p\}$ and a subsequence $p_{n}$ of $\mathbb{N}$ such that
$d(\mathbf{a}_{i_{0}},\mathbf{b}_{r_{p_{n}}}^{p_{n}})<1/p_{n}.$
Similarly, there exists $r_{0}\in\{1,\ldots,Q\}$ and a subsequence $(k_{n})$ of $(p_{n})$ such that
$d(\mathbf{a}_{i_{0}},\mathbf{b}_{r_{0}}^{k_{n}})<1/k_{n}.$
Now we have
$d\left(\mathbf{a}_{j_{k_{n}}},{(f^{\times N})}^{m_{k_{n}}}(\mathbf{b}_{l_{k_{n}}}^{k_{n}})\right)<1/k_{n}.$
By repeating the same reasoning once again, we find a new subsequence of $(k_{n})$, denote it with $(s_{n})$, and $j_{0}\in\{1,\ldots,m\}$ such that
$d\left(\mathbf{a}_{j_{0}},{(f^{\times N})}^{m_{s_{n}}}(\mathbf{b}_{l_{s_{n}}}^{s_{n}})\right)<1/s_{n}.$
Furthermore, as before, there exists $l_{0}\in\{1,\ldots,Q\}$ and a subsequence $(t_{n})$ of $(s_{n})$ such that
$d\left(\mathbf{a}_{j_{0}},{(f^{\times N})}^{m_{t_{n}}}(\mathbf{b}_{l_{0}}^{t_{n}})\right)<1/t_{n}.$
To summarize, we have
$d(\mathbf{a}_{i_{0}},\mathbf{b}_{r_{0}}^{t_{n}})<1/t_{n},\quad d\left(\mathbf{a}_{j_{0}},{(f^{\times N})}^{m_{t_{n}}}(\mathbf{b}_{l_{0}}^{t_{n}})\right)<1/t_{n}.$
If we denote $\mathbf{a}_{i_{0}}=(\alpha_{1},\ldots,\alpha_{N})$ and $\mathbf{b}_{r_{0}}^{t_{n}}=(\beta^{t_{n}}_{1},\ldots,\beta^{t_{n}}_{N})$, then $d(\alpha_{j},\beta^{t_{n}}_{j})<1/t_{n}$, for all $1\leqslant j\leqslant N$. Similarly, if we denote $\mathbf{a}_{j_{0}}=(\gamma_{1},\ldots,\gamma_{N})$ and $\mathbf{b}_{l_{0}}^{t_{n}}=(\delta^{t_{n}}_{1},\ldots,\delta^{t_{n}}_{N})$, we have that $d(\gamma_{j},f^{m_{t_{n}}}(\delta^{t_{n}}_{j}))<1/t_{n}$, for all $1\leqslant j\leqslant N$.
NN-tuples $\mathbf{a}_{i_{0}}$ and $\mathbf{a}_{j_{0}}$ both have exactly $k$ mutually different coordinates. We discard all the sequences $(\beta^{t_{n}}_{j})$, and alike $f^{m_{t_{n}}}(\delta^{t_{n}}_{j})$, that have the same limit, and leave just one with the said limit. We are left with $k$ sequences $(\beta^{t_{n}}_{j})$ and $k$ sequences $(f^{m_{t_{n}}}(\delta^{t_{n}}_{j}))$. We will now use simpler notation for the obtained inequalities. We have $d(p_{j},c^{n}_{j})<1/t_{n}$ and $d(r_{j},f^{u_{n}}(d^{n}_{j}))<1/t_{n}$, for all $1\leqslant j\leqslant k$. Note that $A=\{a_{1},\ldots,a_{k}\}=\{p_{1},\ldots,p_{k}\}=\{r_{1},\ldots,r_{k}\}$ and $\{c^{n}_{1},\ldots,c^{n}_{k}\}=\{d^{n}_{1},\ldots,d^{n}_{k}\}$. As before, denote $\mathbf{p}=(p_{1},\ldots,p_{k})$, $\mathbf{r}=(r_{1},\ldots,r_{k})$, $\mathbf{c}^{n}=(c^{n}_{1},\ldots,c^{n}_{k})$ and $\mathbf{d}^{n}=(d^{n}_{1},\ldots,d^{n}_{k})$.
Let $\varepsilon>0$ and choose $\delta<\varepsilon/2$ such that
$d(x,y)<\delta\Rightarrow d(f^{\times k}(x),f^{\times k}(y))<\varepsilon$ (3)
and $n_{0}\in\mathbb{N}$ such that
$d(\mathbf{p},\mathbf{c}^{n})<\delta,\quad d(\mathbf{r},(f^{\times k})^{u_{n}}(\mathbf{d}^{n}))<\varepsilon/2$
for $n\geqslant n_{0}$. This also implies that for every $s\in\{0,\ldots,l-1\}$ and $n\geqslant n_{0}$
$d(\mu^{s}(\mathbf{r}),(f^{\times k})^{u_{n}}(\mu^{s}(\mathbf{d}^{n})))<\varepsilon/2,$ (4)
since the distance doesn’t change if we permute both $k$-tuples in the same way.
Let $\nu$ be the permutation in $S_{k}$ mapping $(c^{n_{0}}_{1},\ldots,c^{n_{0}}_{k})=\mathbf{c}^{n_{0}}$ to $(d^{n_{0}}_{1},\ldots,d^{n_{0}}_{k})=\mathbf{d}^{n_{0}}$ and $m$ such that $\nu^{m}=Id$. Denote $\mathbf{p}_{1}=\nu(\mathbf{p})$. Let $\mu$ be the permutation in $S_{k}$ mapping $\mathbf{p}_{1}$ to $(r_{1},\ldots,r_{k})=\mathbf{r}$ and $l$ such that $\mu^{l}=Id$.
We have that
$d(\mathbf{p},\mathbf{c}^{n_{0}})<\delta\Leftrightarrow d(\nu(\mathbf{p}),\nu(\mathbf{c}^{n_{0}}))<\delta\Leftrightarrow d(\mathbf{p}_{1},\mathbf{d}^{n_{0}})<\delta.$ (5)
[…]
$d(f^{\times k}(\mathbf{p}_{1}),f^{\times k}(\mathbf{d}^{n_{0}}))<\varepsilon.$ (6)
Furthermore,
$d(\mu(\mathbf{d}^{n_{0}}),(f^{\times k})^{u_{n_{0}}}(\mathbf{d}^{n_{0}}))$ $\leqslant d(\mu(\mathbf{d}^{n_{0}}),\mathbf{r})+d(\mathbf{r},(f^{\times k})^{u_{n_{0}}}(\mathbf{d}^{n_{0}}))$
$=d(\mathbf{d}^{n_{0}},\mu^{-1}\mathbf{r})+d(\mathbf{r},(f^{\times k})^{u_{n_{0}}}(\mathbf{d}^{n_{0}}))$
$=d(\mathbf{d}^{n_{0}},\mathbf{p}_{1})+d(\mathbf{r},(f^{\times k})^{u_{n_{0}}}(\mathbf{d}^{n_{0}}))$
$<\varepsilon/2+\varepsilon/2=\varepsilon,$
Therefore, for every $s\in\{0,\ldots,l-1\}$, we have that
$d(\mu^{s}(\mathbf{d}^{n_{0}}),(f^{\times k})^{u_{n_{0}}}(\mu^{s-1}(\mathbf{d}^{n_{0}})))<\varepsilon.$ (7)
[…]
$\mathbf{p}_{1},$
$f^{\times k}(\mathbf{d}^{n_{0}}),(f^{\times k})^{2}(\mathbf{d}^{n_{0}}),\ldots,(f^{\times k})^{u_{n_{0}}-1}(\mathbf{d}^{n_{0}}),$
$\mu(\mathbf{d}^{n_{0}}),f^{\times k}(\mu(\mathbf{d}^{n_{0}})),(f^{\times k})^{2}(\mu(\mathbf{d}^{n_{0}})),\ldots,(f^{\times k})^{u_{n_{0}}-1}(\mu(\mathbf{d}^{n_{0}})),$
$\mu^{2}(\mathbf{d}^{n_{0}}),f^{\times k}(\mu^{2}(\mathbf{d}^{n_{0}})),(f^{\times k})^{2}(\mu^{2}(\mathbf{d}^{n_{0}})),\ldots,(f^{\times k})^{u_{n_{0}}-1}(\mu^{2}(\mathbf{d}^{n_{0}})),$
$\ldots$
$\mu^{l-1}(\mathbf{d}^{n_{0}}),f^{\times k}(\mu^{l-1}(\mathbf{d}^{n_{0}})),(f^{\times k})^{2}(\mu^{l-1}(\mathbf{d}^{n_{0}})),\ldots,(f^{\times k})^{u_{n_{0}}-1}(\mu^{l-1}(\mathbf{d}^{n_{0}})),$
$\mu^{l-1}(\mathbf{r})=\mathbf{p}_{1}$
@@ 4 Main results and proofs
Let $f:X\to X$ be a homeomorphism with finitely many chain recurrent points. Since $\mathcal{CR}(f)$ is finite and invariant, we conclude it consists of only periodic points. We can pass from $f$ to $f^{k}$, for some $k\geqslant 1$, so that all the chain recurrent points become fixed. This change does not affect the polynomial entropy, since we have that $h_{pol}(f^{k})=h_{pol}(f)$. It is clear that this is also true for the map $F_{n}(f)$, i.e.
@@
$=\left\{\begin{array}[]{ll}qF_{n}(f)^{k}q^{-1}(\mathcal{A})&\text{if }\mathcal{A}\neq F_{X}\\ F_{X}&\text{if }\mathcal{A}=F_{X}\end{array}\right$.
@@
so we can, from now on, consider homeomorphisms $f$ with $\mathcal{CR}(f)=Fix(f)$.
@@
Let $X$ be a compact space and $f:X\to X$ a homeomorphism with a finite chain recurrent set $\mathcal{CR}(f)=Fix(f)$. Then $\mathcal{S}F_{n}(f))$ has a finite non-wandering set and $NW(\mathcal{S}F_{n}(f)))=Fix(\mathcal{S}F_{n}(f)))$, for all $n\geqslant 2$.
@@
Since $U$ was an arbitrary neighbourhood of $A$, we see that $A\in NW(F_{n}(X))$. From Theorem 7, we have that $NW(F_{n}(X))=Fix(F_{n}(X))$, and finally:
@@
Let $X$ be a compact space, $f:X\to X$ a homeomorphism with a finite chain recurrent set and $n\geqslant 2$. Then
@@
As we explained, we can consider all the chain recurrent (and then consequently non-wandering) points to be fixed points. Let
@@
First note that $\pi^{-1}(G_{n})$ is $f^{\times n}$-invariant, since $f$ and the restriction of $\pi$ are bijections. Following the same reasoning as before, there is either exactly one non-wandering point, or there is none (either instantaneously, or after removing all the fixed points), so we add one point to the sets $\pi^{-1}(G_{n})$ and $G_{n}$ - point (x~,,x~)($\tilde$$\ldots,\tilde{x}$) and $\{\tilde{x}\}$, respectively. $H$ will remain a homeomorphism, and the entropies are not changed. We can conclude that
@@
Let $X$ be a compact space, $f:X\to X$ a homeomorphism with a finite chain recurrent set and $n>m\geqslant 1$. Then
@@
As we mentioned in the introduction, we obtained the results regarding the polynomial entropy on $F_{n}(X)$, for specific spaces $X=[0,1]$ and $\mathbb{S}^{1}$ in [12]. We complement them with the results on the symmetric product suspension, using the main Theorem 9 and the fact that $h_{pol}(f,[0,1])=h_{pol}(f,\mathbb{S}^{1})=1$ when the set $NW(f)$ is finite. Again, since $[0,1]$ and $\mathbb{S}^{1}$ are regular curves, in this case the non-wandering set and the chain recurrent set coincide.
@@
We will use the same notation as in the proof of Theorem 6 in [13]. Define $x_{n}=(f^{n}(x_{0}))$, $n\in\mathbb{Z}$ and
@@
We can assume that $f$ is strictly increasing and that $f\neq\mathrm{Id}$ (otherwise, consider $f^{2}$ instead of $f$). Since in this case Fix(f)=NW(f){Fix}(f)={NW}(f), there exists at least one wandering point $x_{0}$. In order to deduce this result we only have to comment on the fact that $\alpha_{f}(x_{0})$ and $\omega_{f}(x_{0})$ are in fact one-element sets, since $\lim\limits_{n\to\pm\infty}f^{n}(x_{0})$ exists. ∎
@@
Acknowledgements. The author is very grateful to Jelena Katić for many constructive discussions regarding non-wandering and chain-recurrent sets on hyperspaces.